OpenStudy (anonymous):

Please! How do I simplify? Help me! sen²x/1-sen²x * cotg² x

OpenStudy (wolf1728):

Is sen secant?

OpenStudy (anonymous):

No. Seno.

OpenStudy (wolf1728):

What is seno?

OpenStudy (anonymous):

I think you call sine. Sorry

OpenStudy (wolf1728):

Oh okay

OpenStudy (wolf1728):

sine² x / (1-sine² x) * cotangent² x

OpenStudy (anonymous):

yes

OpenStudy (wolf1728):

sine = opp / hyp cotangent = adj / opp 

OpenStudy (anonymous):

Yes! Please! help me simplify this :(

OpenStudy (wolf1728):

I've looked at it for a long time and I don't have any ideas. Sorry.

OpenStudy (anonymous):

ok.. thanks :)

OpenStudy (unklerhaukus):

sen = Seno = Sine = sin cotg = ? = cotangent = cot = 1/tan = 1/(sin/cos) = cos/sin \[\frac{\sin^2x}{1-\sin^2x} \times \cot^2 x\] use \[\sin^2x+\cos^2x=1\]\[\cos^2x=1-\sin^2x\] to simplify the denominator then use the definition of cotangent to cancel some factors